OpenStudy (anonymous):

Solve the equation x-5/4 = 3 for x. 17 7 12 10

OpenStudy (anonymous):

x = 5/4 +3

mathslover (mathslover):

\[\huge{x-\frac{5}{4}=3}\] \[\huge{x-\frac{5}{4}+\frac{5}{4}=3+\frac{5}{4}}\] \[\huge{x=3+\frac{5}{4}}\]

mathslover (mathslover):

Can you solve this ? @rox13kh u r requested not to answer the questions directly . Thanks !

OpenStudy (anonymous):

sorry T___T